OpenStudy (anonymous):

I'm still having trouble writing these in a short "sum" form. I have the solution, but for whatever reason I'm not able to reason my way to the solution. \[n=0, a_2=\frac{a_0}{2!}\] \[n=1, a_3=\frac{a_1}{3!}\] \[n=2, a_4=\frac{a_0}{4!}\] \[n=3, a_5=\frac{a_1}{5!}\] \[n=4, a_6=\frac{a_0}{6!}\] \[n=5, a_7=\frac{a_1}{7!}\] \[n=6, a_8=\frac{a_0}{8!}\] I understand that I will have \[y=a_0\sum +a_1\sum\]

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@blues

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

Originally I had \[a_{n+2}=\frac{a_n}{(n+1)(n+2)}\]

OpenStudy (zarkon):

are you trying to write \[y=\sum_{n=0}^{\infty}a_n\] but then to break it up into two sums?

OpenStudy (anonymous):

If one sum is sufficient we can write it in one sum, but how do I know that I need to sums or one?

OpenStudy (zarkon):

two sums would be nice since they would be in terms of \(a_0\) and \(a_1\)

OpenStudy (zarkon):

if you look t the even terms... \[a_0+a_2+a_4+a_6+\cdots\] \[a_0+\frac{a_0}{2!}+\frac{a_0}{4!}+\frac{a_0}{6!}+\cdots\] \[a_0\left(1+\frac{1}{2!}+\frac{1}{4!}+\frac{1}{6!}+\cdots\right)\] \[a_0\left(\frac{1}{0!}+\frac{1}{2!}+\frac{1}{4!}+\frac{1}{6!}+\cdots\right)\] \[a_0\left(\frac{1}{0!}+\frac{1}{2!}+\frac{1}{4!}+\frac{1}{6!}+\cdots\right)=a_0\sum_{k=0}^{\infty}\frac{1}{(2k)!}\]

OpenStudy (anonymous):

makes sense! Let's see if I can do a_1

OpenStudy (anonymous):

\[a_3+a_5+a_7+a_9\] \[\frac{a_1}{3!}+\frac{a_1}{5!}+\frac{a_1}{7!}+\frac{a_1}{9!}+\dots\] \[a_1\left(\frac{1}{3!}+\frac{1}{5!}+\frac{1}{7!}+\frac{1}{9!}+\dots\right)=a_1\sum_{k=0}^{\infty}\frac{1}{2k+1}\]

OpenStudy (anonymous):

ooops!

OpenStudy (anonymous):

\[a_1\left(\frac{1}{3!}+\frac{1}{5!}+\frac{1}{7!}+\frac{1}{9!}+\dots\right)=a_1\sum_{k=0}^{\infty}\frac{1}{(2k+1)!}\]

OpenStudy (zarkon):

right...we should probably start both sums at k=1

OpenStudy (anonymous):

sure

OpenStudy (zarkon):

since looking above your n=0 corresponds to \(a_2\)

OpenStudy (zarkon):

\[y=a_0\sum_{k=1}^{\infty}\frac{1}{2k!}+a_1\sum_{k=1}^{\infty}\frac{1}{(2k+1)!}\]

OpenStudy (zarkon):

\[y=a_0\sum_{k=1}^{\infty}\frac{1}{(2k)!}+a_1\sum_{k=1}^{\infty}\frac{1}{(2k+1)!}\]

OpenStudy (anonymous):

Finally It makes sense. I think writing out the final step with the big parenthesis allowed me to see the sum. THanks a million!

OpenStudy (zarkon):

np